OpenStudy (anonymous):

What is the 43rd term of an arithmetic sequence with a rate of increase of -6 and a11 = 12? A)-174 B)-176 C)-180 D)-186 E)-240

OpenStudy (anonymous):

@caleb_gajewsky

OpenStudy (anonymous):

_>_<_

OpenStudy (anonymous):

I don't understand the "a11 = 12" part :/

OpenStudy (anonymous):

\[a ^{11} =12\]

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

sorry qt

OpenStudy (dan815):

\[a_11=12 \\ a_12=12-6=6\\a_{13}=6-6=0 \\ a_{14}=0-6=-6\\ a_{15}=.....\\.\\.\\.\\.\\ a_{43}=?\]

OpenStudy (dan815):

how many times will we hhave to subtract 6? to get from term 11 to term 43

OpenStudy (dan815):

Think about that and get back to me with the answer :)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

omg idk and im stuck on a test todays the last day please give me the answers i know normally you shldnt and normally i dont ask for just answers but i need them now omg D:

OpenStudy (dan815):

okay lets see how many numbers are betwen 11 to 43

OpenStudy (dan815):

*summer* T_T

OpenStudy (dan815):

pay attention then

OpenStudy (anonymous):

omg im about to fail

OpenStudy (anonymous):

-186?

OpenStudy (dan815):

we are subtracting from the 12th term so... 12 to 43 is how many numbers if i subtract 11 from 12 and 11 from 43, that wont change the total numbers i have 12-11 to 43 -11 1 to 32 <---- so there are a total of 32 numbers between from 12 to 43

OpenStudy (anonymous):

i have 1 more minute to answer this question

OpenStudy (dan815):

we are subtracting 32 times from 12 12-(32*6)=-180

OpenStudy (anonymous):

it logged me out before i could finish omg D;

OpenStudy (anonymous):

but thanx i know how to do it now

OpenStudy (dan815):

:) dont feel bad

OpenStudy (dan815):

*hug*

OpenStudy (dan815):

*runs his fingers through his beautiful hair*

OpenStudy (anonymous):

lol